OpenStudy (anonymous):

Decide whether or not cos^2 x(1+tan^2 x) = 1. Explain your reasoning

OpenStudy (anonymous):

\[Cos^2(x) * (1+\tan^2(x))\] Is this your question?

OpenStudy (mathstudent55):

\(cos^2 x(1+tan^2 x) = 1\) Let's use this identity (in red, below): \(\color {red}{\sin^2 x + cos^2 x = 1}\) Divide both sides by \( \cos^2 x\). \(\color{red}{\tan^2 x + 1 = \sec^2 x}\) Solve for \(\tan^2 x\). \(\color{red}{\tan^2 x = \sec^2 x - 1}\) Now substitute what \(\tan^2 x \) is equal to in the original equation. \(cos^2 x(1+\color{red}{\sec^2 x - 1}) = 1\) Can you continue from here?

OpenStudy (anonymous):

I'm still not sure. What am I looking for to see if it is an identity or not?

OpenStudy (mathstudent55):

To prove that an equation is an identity, you must show that the left side and right side equal each other.

OpenStudy (anonymous):

so I have to make the left side = 1?

OpenStudy (mathstudent55):

Correct. We're almost there. Subtract the 1 and 1 in the parentheses. What do you have now?

OpenStudy (anonymous):

cos^2 x(sec^2 x)=1

OpenStudy (mathstudent55):

What is \( \sec^2 x\) equal to? (in terms of \( \cos^2 x\))

OpenStudy (anonymous):

hmm not sure. tan^2x

OpenStudy (mathstudent55):

\( \sec^2 x = \dfrac{1}{\cos^2 x}\)

OpenStudy (anonymous):

ok, so it is a trig identity right?

OpenStudy (mathstudent55):

Correct, because the left side equals 1.

OpenStudy (anonymous):

Ah ok so to clarify: I would just explain the reasoning by writing both sides have to be equal and show that sec^2 x = 1/cos^2x

OpenStudy (anonymous):

I just put the equation into symbolab and it said the identity was false? @mathstudent55

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Who is "it"?

OpenStudy (anonymous):

Hmm that's weird then! I was just confused because the program said it was false. Thanks for explaining and linking me to Wolf

hero (hero):

1. Given: \[\cos^2 x(1+\tan^2 x) = 1\] 2. Re-write \(\tan^2x(x)\) as \(\dfrac{\sin^2x}{\cos^2x}\): \[\cos^2x\left(1 + \frac{\sin^2x}{\cos^2x}\right) = 1\] 3. Factor out \(\dfrac{1}{\cos^2x}\): \[\frac{\cos^2x}{\cos^2x} \left( \cos^2x + \sin^2x \right) = 1\] 4. From there, it is easy to see that: \[1(1) = 1\]